S1. The area bounded by the curves y=x^2-3 and y=kx+2 is the least,if k=0.


S2:The area bounded by the curves y=x^2-3 and y=kx+2 is (k^2+20)^1/2.


a. is true;s2 is a correct explanation for S1.


b.S1 is true ;S2 is not a correct explanation for S1.


c.S1 is true,S2 is false


d.S1 is false,S2 is true

Grade:12

1 Answers

Ramesh V
14 years ago

The answer is : c.)S1 is true,S2 is false

The area bounded by the curves y=x2- 3 and y=2( for k=0 in y=kx+2)  is

2*integral of (y+3)1/2dy; limits from -3 to 2

area is 20*51/2/3

 

so for K=0 putting k=0 in S2 the area bounded does not come the same from formula (k^2+20)1/2 as in S2

 

So S1 is true,S2 is false